OpenStudy (anonymous):

if A ,P,D are square matrix and A=P*D*P^-1 prove that a)A^2=P*D^2*P^-1 b)prove with induction for every m>=0 we have A^m=P*D^m*P^-1

OpenStudy (jamesj):

Have you tried calculating? Part a is very easy. Just do the calculation.

OpenStudy (anonymous):

yes IS IT this way A^2 =P*P*D^2*P^-1*P^-1 SINCE P*P^-1=I A^2=P*D^2*P^-1?

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

can you give me a little help for part b) i know how to prove it for m=1 for m=kbut i don't know hoe to continue

OpenStudy (anonymous):

how*

OpenStudy (jamesj):

Let P(j) be the hypothesis that \[ A^j = PD^jP^{-1} \] Now P(0) is true, and you also know that P(1) and part (a) shows P(2) is true.

OpenStudy (jamesj):

Now you need to show that if P(k) is true, that this implies P(k+1).

OpenStudy (jamesj):

Suppose then P(k). That is, \[ A^k = PD^kP^{-1} \] Now \[A^{k+1} = A^k A \] \[ = (PD^kP^{-1})A, \ \ \ \ \hbox{by hypothesis} \] \[ = ... \] you take it from here.

OpenStudy (anonymous):

in the end it is A^k+1=P*P^-1*D^k+1

OpenStudy (anonymous):

can you give me an advice for this exercise too if C is a square matrix that proves this C^2+C=I then PROVE THAT C^4=2I-3C

OpenStudy (jamesj):

\[ (PD^kP^{-1})A \] \[= (PD^kP^{-1})(PDP) \] \[ = PD^k(P^{-1}P)DP^{-1} \] \[ = PD^kIDP^{-1} \] \[ = PD^{k+1}P^{-1} \]

OpenStudy (anonymous):

i did it this way